Fosnight Enterprises prepared the following sales budget Month March April Budgeted Sales $9,000 $14,000 O A. $11,280 OB. $11,120 ?. $12.320 O D. $1.120 Click to select your answer ch
media%2F430%2F4302248b-fcc9-489a-b3ae-f1
media%2Fc41%2Fc419bdc1-949f-4358-a066-f7
Answer #1

B. $11,120

Cost of Goods Sold = $14,000 × 80% = $11,200

Ending Inventory = $13,000 × 80% × 10% = $1,040

Beginning Inventory = $14,000 × 80% × 10% = $1,120

Purchases = $11,200 + $1,040 - $1,120 = $11,120